# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v1

@@ I Introduction
Managing energy assets in a grid system is a complex task that requires decision-making under uncertainty. The dynamics and stochasticity of the environment create a constantly evolving and inherently complex system. This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization. In many applications, the future is reasonably predictable. Indeed, it is possible to make predictions about future cost function with a significant accuracy, particularly near future.
@@ I-A Related work
The problem of decision making with no complete information is often called an online optimization problem. According to Powell [10], the problem involves two main challenges: modelling the sequence of decisions and designing a policy to maximize the objective. Our focus is on a type of optimization problem that depends on the current state, uses a direct lookahead approach, and includes switching costs. The exact details of this problem are given in the Problem Statement in the Section II-A.
@@
Current work sets to explore the performance of commitment-based policies in the context of energy management systems. We are particularly interested in analyzing the forecast and planning stability, as a potential factor that influences the performance of the policy. A comprehensive study of forecast stability is made by Godahewa et al. [18]. Those authors categorize the stability into two types: vertical and horizontal. Vertical stability refers to the variation period-to-period variability between forecasts given for the same time period but issued from different origins. Horizontal stability describes the variation between predictions made at one instance within a prediction window. The lack of planning stability has long been recognized an important issue within research on supply chain management. The vertical stability in order forecasts is often referred to as the system nervousness. It is a well-known phenomenon that can lead to inefficient production and inventory management. The extra capacity implications are also referred to as the bullwhip effect [19]. The commonly used methods to obtain a stable plan include freezing intervals [20] [21], using smoothing techniques [22], and incorporating changing costs into the objective function [23] [24].
Freezing order plans is synonymous with the concept of commitment, discussed earlier in the context of OCO problems with switching costs. The freezing intervals are periods during which the plan is not revised, and the production is executed according to the plan. In supply chain management, the freezing is found to be effective in reducing the nervousness of the system [20]. However, it is also found that the freezing itself is a potential source of nervousness. It is shown by Tang and Grubbström [25] that forecast errors have an effect on the selection of optimal freezing intervals. Xie et al. [26] show that the interaction between forecast errors and the freezing intervals has a significant impact on the performance of the system.
@@ I-B Contents and Contributions
  • First, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
@@ II-A Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [32] and based on a real-world dataset from a grid-interactive community. More generally, we consider an online optimization with switching costs. The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Appendix A-A. The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [10], we use the following formulation:
@@
$+\beta\left\lVert\tilde{x}_{tt^{\prime}}-\tilde{x}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
@@
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Appendix A-A.
@@ II-B Optimization
II-C Theoretical Analysis of Performance Bounds
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (4)
Independent and Identically Distributed Noise
$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
The statistical properties of prediction errors in this context are:
$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
$\operatorname{trace}(R_{e})=\sigma^{2}$
This leads to the upper bound on the FHC cost:
Exponentially Decaying Correlation
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (7)
We then express the prediction error covariance as:
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (8)
$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (9)
To simplify, we group the parameters independent of $v$:
$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
Thus, the expected competitive difference of the FHC algorithm is:
Stochastic Optimization
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (12)
II-D Forecast
@@
While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings. This setup follows the problem setting given in the Citylearn Challenge 2022 [32]. Although this train-test splitting is not typical in forecasting, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
II-D1 Rolling Origin Window
@@ II-D2 Stability
Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC). The formulations of the stability metrics vary for vertical and horizontal types. Following the general framework of Van Belle et al. [36], we evaluate stability for deterministic forecasts. In line with the definition of sMAPC in [36], we define the absolute scale stability metric, Mean Absolute Change (MAC). The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
@@
II-D3 Stability of Probabilistic Forecasts
The Earth Mover’s Distance (EMD), also known as the Wasserstein distance, quantifies the minimum amount of “work” required to transform one distribution into another, making it particularly suitable for assessing the stability of scenario sets. In a stochastic optimization problem, the choice of EMD for measuring the stability of probabilistic forecasts is based on its advantageous properties for comparing empirical distributions. Firstly, EMD is a symmetric measure and holds a triangle inequality, which makes it a valid metric. Secondly, it does not require both distributions to be on the same probability space, which makes it effective even when the two sets of forecasts do not overlap—for instance, when an initial forecast predicts a sunny day and high photovoltaic generation, and an updated forecast adjusts the prediction to a cloudy day and low generation. The EMD provides insights into the extent of the adjustments between the old predictions and the new ones. Lastly, EMD takes into account the dimension of the metric space, which makes it equivalent to the mean absolute change (MAC) used for point forecasts.
Several other metrics, such as Kullback-Leibler (KL) divergence and Jensen-Shannon distance, can measure statistical distance between distributions. However, EMD is particularly suitable for comparing scenario sample sets due to the reasons outlined above. Formally, in a 1-dimensional case, where $P$ and $Q$ are empirical sets of samples $P_{1},...,P_{n}$ and $Q_{1},...,Q_{m}$, the EMD is defined as:
$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|X_{(i)}-Y_{(i)}\right\|$ (15)
where $X_{(i)}$ and $Y_{(i)}$ are the sorted samples of $P$ and $Q$ respectively.
This metric measures the dissimilarity between two sets of scenarios. To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability. Horizontal stability (${EMD}_{H}$) measures the similarity between predictions within the same horizon, while vertical stability (${EMD}_{V}$) evaluates the distance between predictions generated for the same period at different origins:
@@ II-D3 Stability of Probabilistic Forecasts
II-D4 Accuracy of Determinstic Forecasts
@@
II-D5 Accuracy of Probabilistic Forecasts
@@ III Experimentation and Results
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in Appendix B. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
@@
Figure 4: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from out-of-sample buildings. The optimization score is shown with and without switching costs. The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD. The optimization score is measured with the cost function of the FHC algorithm.
@@ IV Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@ A-B MPC Formulation
Appendix B In-sample Performance

### v2

@@ I Introduction
Managing energy assets in a grid system is a complex task that requires decision-making under uncertainty. The dynamics and stochasticity of the environment create a constantly evolving and inherently complex system. This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization.
@@ I-A Related work
Current work sets to explore the performance of commitment-based policies in the context of energy management systems. We are particularly interested in analyzing the forecast and planning stability, as a potential factor that influences the performance of the policy. A comprehensive study of forecast stability is made by Godahewa et al. [17]. Those authors categorize the stability into two types: vertical and horizontal. Vertical stability refers to the variation period-to-period variability between forecasts given for the same time period but issued from different origins. Horizontal stability describes the variation between predictions made at one instance within a prediction window. The lack of planning stability has long been recognized an important issue within research on supply chain management. The vertical stability in order forecasts is often referred to as the system nervousness. It is a well-known phenomenon that can lead to inefficient production and inventory management. The extra capacity implications are also referred to as the bullwhip effect [18]. The commonly used methods to obtain a stable plan include increased commitment, referred as freezing intervals [19] [20], using smoothing techniques [21], and incorporating changing costs into the objective function [22] [23].
@@ I-B Contents and Contributions
  • First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs. We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
@@ II-A Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [29] and based on a real-world dataset from a grid-interactive community. More generally, we consider an online optimization with switching costs. The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Online Appendix A [^1] Appendix A. [Online] Available: https://arxiv.org/abs/2407.03368. The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [30], we use the following formulation:
@@
$+\beta\left\lVert\tilde{y}_{tt^{\prime}}-\tilde{y}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
@@
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Online Appendix A  [^2] Appendix A. [Online] Available: https://arxiv.org/abs/2407.03368. Furthermore, In the Online Appendix C  [^3] Appendix C. [Online] Available: https://arxiv.org/abs/2407.03368, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
@@ II-B Optimization
II-C Forecast
@@
While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings. This setup follows the problem setting given in the Citylearn Challenge 2022 [29]. Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
II-C1 Rolling Origin Window
@@ II-C2 Stability
Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC). The formulations of the stability metrics vary for vertical and horizontal types. In line with the definition of sMAPC in [34], we define the absolute scale stability metric, Mean Absolute Change (MAC). The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
@@
II-C3 Stability of Probabilistic Forecasts
The Earth Mover’s Distance (EMD), or Wasserstein distance, quantifies the minimum ”work” required to transform one distribution into another, making it ideal for assessing the stability of scenario sets in stochastic optimization. EMD is chosen for its symmetric nature, adherence to the triangle inequality, and ability to compare distributions on different probability spaces, making it effective even when forecasts differ significantly. It also accounts for the metric space dimension, akin to the mean absolute change (MAC) used for point forecasts. Other metrics like Kullback-Leibler (KL) divergence and Jensen-Shannon distance measure statistical distance, but EMD is preferred for scenario sample sets. In a 1-dimensional case with empirical sets $P$ and $Q$, EMD is defined as:
$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|P_{(i)}-Q_{(i)}\right\|$ (6)
where $P_{(i)}$ and $Q_{(i)}$ are the sorted samples of $P$ and $Q$ respectively. This metric measures the dissimilarity between two sets of scenarios. To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability:
@@ II-C3 Stability of Probabilistic Forecasts
II-C4 Accuracy of Determinstic Forecasts
@@
II-C5 Accuracy of Probabilistic Forecasts
@@ III Experimentation and Results
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in Online Appendix D  [^4] Appendix D. [Online] Available: https://arxiv.org/abs/2407.03368. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
@@
Figure 4: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings. The optimization score is shown with and without switching costs. The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD. The optimization score is measured with the cost function of the FHC algorithm.
@@ IV Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis in Online Appendix C, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@ A-B MPC Formulation
A-C Theoretical Analysis of Performance Bounds
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (27)
Independent and Identically Distributed Noise
$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
The statistical properties of prediction errors in this context are:
$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
$\operatorname{trace}(R_{e})=\sigma^{2}$
This leads to the upper bound on the FHC cost:
Exponentially Decaying Correlation
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (30)
We then express the prediction error covariance as:
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (31)
$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (32)
To simplify, we group the parameters independent of $v$:
$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
Thus, the expected competitive difference of the FHC algorithm is:
Stochastic Optimization
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (35)
A-D In-sample Performance